### v2

# Eigenvalue bounds for Schrödinger operators with complex potentials on compact manifolds

Abstract.
We prove eigenvalue bounds for Schrödinger operator $\Delta_{g}+V$ on compact manifolds with complex potentials $V$. The bounds depend only on an $L^{q}$-norm of the potential, and they are shown to be optimal, in a certain sense, on the round sphere and more general Zoll manifolds. These bounds are natural analogues of Frank’s [9] results in the Euclidean case.
2020 Mathematics Subject Classification:
58J50, 35P15, 31Q12
Support through the Engineering & Physical Sciences Research Council (EP/X011488/1) is acknowledged.
1. Introduction and main results
1.1. Spectral inclusion
Let $\Delta_{g}$ be the Laplace-Beltrami operator on a closed (i.e., compact boundaryless) Riemannian manifold $(M,g)$ of dimension $n\geq 2$. We consider Schrödinger operators
(1) $-\Delta_{g}+V\quad\mbox{on }L^{2}(M)$
with a complex-valued potential $V$. The unperturbed operator $\Delta_{g}$ is a non-negative, unbounded self-adjoint operator in $L^{2}(M)$, defined with respect to the Riemannian volume measure ${\rm d}x$. Since $\Delta_{g}$ has compact resolvent, its spectrum consists of nonnegative eigenvalues $0=\lambda_{0}^{2}\leq\lambda_{1}^{2}\leq\ldots$ with $\lambda_{k}^{2}\to\infty$, and the set of corresponding eigenfunctions $(e_{j})$ forms an orthonormal basis of $L^{2}(M)$. By standard perturbation theory, if $V\in L^{\infty}(M)$, then the eigenvalues of $\Delta_{g}+V$ are contained in a union of closed disks,
(2) $\operatorname{spec}(-\Delta_{g}+V)\subset\bigcup_{k=0}^{\infty}D(\lambda_{k}^{2},r),\quad r=\|V\|_{L^{\infty}(M)}.$
In this paper, we deal with unbounded potentials, i.e., $V\in L^{q}(M)$ for some $q<\infty$.
Theorem 1.1.
Let $\Delta_{g}$ be the Laplace-Beltrami operator on a closed Riemannian manifold $(M,g)$ of dimension $n\geq 2$, and let $q>n/2$. Then there exists a constant $C=C(M,g,q)$ such that for all $V\in L^{q}(M)$,
(3) $\operatorname{spec}(-\Delta_{g}+V)\subset\bigcup_{k=0}^{\infty}D(\lambda_{k}^{2},Cr_{k})\cup\{z\in{\mathbb{C}}:|z|^{\frac{1}{2}}(1+|z|)^{-\sigma(q)}\leq C\|V\|_{L^{q}(M)}\},$
where $\lambda_{k}^{2}$ are the eigenvalues of $\Delta_{g}$, $r_{k}:=\|V\|_{L^{q}(M)}(1+\lambda_{k})^{2\sigma(q)}$, and
(4) $\sigma(q):=\begin{cases}\frac{n}{2q}-\frac{1}{2}\quad&\mbox{if }\frac{n}{2}\leq q\leq\frac{n+1}{2},\\ \frac{n-1}{4q}\quad&\mbox{if }\frac{n+1}{2}\leq q\leq\infty.\end{cases}$
If $n\geq 3$ and $q=n/2$, then there exist constants $C,c>0$, depending on $M,g,q$, such that for all $V\in L^{\frac{n}{2}}(M)$ satisfying $\|V\|_{L^{\frac{n}{2}}(M)}\leq c$, we have
(5) $\operatorname{spec}(-\Delta_{g}+V)\subset\bigcup_{k=0}^{\infty}D(\lambda_{k}^{2},Cr_{k}),\quad r_{k}:=\|V\|_{L^{\frac{n}{2}}(M)}(1+\lambda_{k}).$
We emphasize that the constant in (3) is independent of $V$. In particular, $V$ is allowed to be frequency-dependent; cf. Theorem 1.3 below. Estimates of a similar type were established in [6] for the Landau Hamiltonian (that is, the Schrödinger operator with constant magnetic field) in ${\mathbb{R}}^{n}$ ($n$ even). In contrast to the situation here, the radii of the disks tend to zero (i.e., the powers of $\lambda_{k}$ are negative) in this case.
Under the assumptions of Theorem 1.1, the operator $\Delta_{g}+V$ can be defined as an $m$-sectorial operator by standard quadratic form methods; see, e.g., [14], or more specifically, [10, Appendix B]. The proof uses Sobolev embedding, as in the case of real-valued $V$, considered in [2, Section 6].
Remark 1.2.
(i) We will refer to the union of disks $\bigcup_{k=0}^{\infty}D(\lambda_{k}^{2},Cr_{k})$ as “the first set” in (3) and to $\{z\in{\mathbb{C}}:|z|^{\frac{1}{2}}(1+|z|)^{-\sigma(q)}\leq C\|V\|_{L^{q}(M)}\}$ as “the second set”. By distinguishing the cases $|z|\leq 1$ or $|z|>1$, it is easy to see that the second set is contained in the disk
(6) $\{z\in{\mathbb{C}}:|z|\leq\max((2^{\sigma(q)}C\|V\|_{L^{q}(M)})^{2},(2^{\sigma(q)}C\|V\|_{L^{q}(M)})^{\frac{1}{1/2-\sigma(q)}})\}$
centered at the origin. Since $\lambda_{0}=0$, the disk $D(\lambda_{0}^{2},Cr_{0})$ is also centered at the origin and could be absorbed in (6) by including the term $C\|V\|_{L^{q}(M)}$ in the maximum. However, the decomposition (3) arises naturally in the proof.
(ii) Asymptotically, as $k\to\infty$, the disks in the first set are contained inside the region
(7) $\{z\in{\mathbb{C}}:|\mathrm{Im}\,z|\leq C\|V\|_{L^{q}(M)}(\mathrm{Re}\,z)^{\sigma(q)}\}.$
Note that $\sigma(q)<1/2$ for $q>n/2$ and $\sigma(q)=1/2$ for $q=n/2$. Hence, for the latter case, the region (7) is the interior of a parabola around the positive $\mathrm{Re}\,z$ axis, while for $q>n/2$, it is much smaller, approaching a strip $|\mathrm{Im}\,z|\leq C\|V\|_{L^{q}(M)}$ as $q\to\infty$.
1.2. Comparison with the Euclidean case
We recall the bounds for the Schrödinger operators on $L^{2}({\mathbb{R}}^{n})$ that we wish to compare our results to: For $n/2<q\leq(n+1)/2$,
(8) $\operatorname{spec}(-\Delta+V)\subset[0,\infty)\cup\Big\{z\in{\mathbb{C}}:|z|^{q-\frac{n}{2}}\leq C_{n,q}\int_{{\mathbb{R}}^{n}}|V(x)|^{q}{\rm d}x\Big\}.$
Moreover, if $n\geq 3$ and $q=n/2$, then there exists a constant $c_{n}$ such that there are no eigenvalues outside $0,\infty)$ if $\|V\|_{L^{n/2}({\mathbb{R}}^{n})}\leq c_{n}$. The one-dimensional bound is due to Abramov–Aslanyan–Davies [1], and the higher dimensional bounds are due to Frank [9]. It is known [3] that the inequality (8) fails for $q>(n+1)/2$, whereas Laptev and Safronov [16] had conjectured its validity up to $q<n$. The limitation to $q\leq(n+1)/2$ essentially comes from the Knapp example in harmonic analysis. It requires that $V$ decay sufficiently fast at infinity in an average sense. On a compact manifold, such a condition is of course meaningless. However, if we consider the $n$-dimensional flat torus $\mathbb{T}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$ and rescale by a factor $L>0$, we obtain
(9) $\operatorname{spec}_{L^{2}(L\mathbb{T}^{n})}(-\Delta+L^{-2}V(\cdot/L))=L^{-2}\operatorname{spec}_{L^{2}(\mathbb{T}^{n})}(-\Delta+V).$
Hence, for the rescaled operator, the decay of the potential is now measured relative to the scale $L$. According to Theorem 1.1, its spectrum is contained in the $L^{-2}$ dilation of the set on the right-hand side of (3). In Section 4, we will show that, as $L\to\infty$, all limit points of this set lie in the right-hand side of (8), provided that $n/2<q\leq(n+1)/2$. More precisely, it follows from the proof of Lemma 4.1 that limit points of the first set in (3) (the union of disks) are contained in $0,\infty)$, while limit points of the second set of (3) are contained in the second set on the right-hand side of (8).
We could also use (3) to prove (8), but since the latter is already known, we will not pursue this direction. More importantly, the argument shows that since (8) is known to be optimal [3], the second set on the right-hand side of (3) is also optimal in the sense that, for large $|z|$, the power $|z|^{\frac{1}{2}-\sigma(q)}$ cannot be improved on a torus, in the regime where the sidelength $L$ tends to infinity.
1.3. Saturation on the sphere and Zoll manifolds
We turn to the optimality of the first set on the right-hand side of (3) when $M=S^{n}$, the round sphere, or more generally, a Zoll manifold. These are manifolds whose geodesics are periodic with a common minimal period. Without loss of generality after rescaling the metric, we may assume that the common minimal period is $2\pi$. By a theorem of Weinstein [23], there is a constant $\alpha=\alpha_{M}\geq 0$ so that all the nonzero eigenvalues of $\Delta_{g}$ cluster around the values $(k+\alpha)^{2}$, $k\in{\mathbb{N}}$. More precisely, there is a constant $A=A_{M}$ such that each non-zero eigenvalue is contained in a cluster $(k+\alpha-A/k)^{2},(k+\alpha+A/k)^{2}]$ for some $k\in{\mathbb{N}}$, and the number of eigenvalues in this cluster, counted with multiplicity, is $d_{k}\approx k^{n-1}$. We label the eigenvalues in the $k$-th cluster by $\lambda_{k,j}^{2}$, $j=1,\ldots,d_{k}$.
In the case of the sphere $S^{n}$, the eigenvalues of the Laplacian are given explicitly by $\lambda_{k,j}^{2}=k(k+n-1)$, and thus the first set in (3) is a union of disks of the form
(10) $D_{k}:=D(k(k+n-1),Ck^{2\sigma(q)}\|V\|_{L^{q}(S^{n})}),$
when $k\geq 1$. We will consider the asymptotic regime $k\to\infty$ here and allow $V$ to depend on $k$. For general Zoll manifolds, we will additionally assume that
(11) $k^{2\sigma(q)}\|V\|_{L^{q}(M)}\gtrsim 1,$
which ensures that all the disks $D(\lambda_{k,j}^{2},Cr_{k,j})$ corresponding to $k$-th cluster are contained in the larger disk
(12) $\tilde{D}_{k}:=D((k+\alpha)^{2},Ck^{2\sigma(q)}\|V\|_{L^{q}(S^{n})}),$
for some larger constant $C$, depending on $A$ and the lower bound in (11).
Theorem 1.3.
i) Let $n\geq 2$ and $q>n/2$. For any $\rho>0$, $\theta\in[0,2\pi)$ and $k\gg 1+\rho^{1/(1-2\sigma(q))}$, there is a potential $V=V_{k,\theta,\rho}\in L^{q}(S^{n})$, $\|V\|_{L^{q}(S^{n})}\approx\rho$, such that
(13) $k(k+n-1)+\rho{\rm e}^{\mathrm{i}\theta}(1+\mathcal{O}(\rho k^{2\sigma(q)-1}))k^{2\sigma(q)}\in\operatorname{spec}(-\Delta_{S^{n}}+V)\cap D_{k}.$
If $n\geq 3$ and $q=n/2$, then (13) holds provided $\rho\ll 1$.
ii) If $M$ is a general Zoll manifold of dimension $n\geq 2$, $q>n/2$, $\rho>0$, $\theta\in[0,2\pi)$ and $k\gg 1$, and if we assume in addition that $k^{2\sigma(q)}\rho\gg 1$, then there is a potential $V=V_{k,\theta,\rho}\in L^{q}(M)$, $\|V\|_{L^{q}(S^{n})}\approx\rho$, such that
(14) $(k+\alpha)^{2}+\rho{\rm e}^{\mathrm{i}\theta}(1+\mathcal{O}(\rho k^{2\sigma(q)-1}))k^{2\sigma(q)}\in\operatorname{spec}(-\Delta_{g}+V)\cap\tilde{D}_{k}.$
If $n\geq 3$ and $q=n/2$, then (14) holds provided $\rho\ll 1$.
Remark 1.4.
(i) The case $q=\infty$ is included but is not particularly interesting, as it is easily seen that (2) is sharp when $V$ is taken to be a constant potential.
(ii) Theorem 1.3 shows that, on the sphere or more general Zoll manifolds, the radius of the disks in Theorem 1.1, $r_{k}=\|V\|_{L^{q}(M)}(1+\lambda_{k})^{2\sigma(q)}$, cannot be decreased (up to the value of the constant). More precisely, (3) cannot hold if $r_{k}$ there is replaced by $\epsilon(\lambda)r_{k}$, where $\epsilon:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ is any decreasing function satisfying $\lim_{\lambda\to\infty}\epsilon(\lambda)=0$.
1.4. Resolvent estimates
Frank [9] proved the Euclidean bound (8) by combining uniform $L^{p}\to L^{p^{\prime}}$ resolvent bounds, established by Kenig, Ruiz, and Sogge [15] in 1987, with the well-known Birman–Schwinger principle. The latter stipulates that $z$ is an eigenvalue of $\Delta+V$ if and only if 1-1 is an eigenvalue of the compact operator $|V|^{1/2}(-\Delta-z)^{-1}V^{1/2}$. Here and in the following, $V^{1/2}:=V/|V|^{1/2}$. We will follow the same strategy for proving Theorem 1.1. Compared to the Euclidean case, $L^{p}\to L^{p^{\prime}}$ resolvent bounds for the Laplace–Beltrami operator on compact manifolds have a comparatively short history. To make a direct connection to Theorem 1.1 we shall adopt the following convention throughout this paper,
(15) $\frac{1}{q}=\frac{1}{p}-\frac{1}{p^{\prime}}.$
Dos Santos Ferreira, Kenig, and Salo [8] proved the uniform bound
(16) $\|(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\lesssim|z|^{\sigma(q)-\frac{1}{2}},\quad|\mathrm{Im}\,\sqrt{z}|\geq 1,$
for $p=2n/(n+2)$ and $n\geq 3$, in which case $\sigma(q)-\frac{1}{2}=0$. The latter improves earlier work of Shen [17] for the flat torus. Here and in the following, we are using the principal branch of the square root on ${\mathbb{C}}\setminus(-\infty,0]$ with $\mathrm{Re}\,\sqrt{z}\geq 0$. Frank and Schimmer [12], and, independently, Burq, Dos Santos Ferreira, and Krupchyk [5], proved (16) for the endpoint $p=2(n+1)/(n+3)$ and $n\geq 2$. A straightforward interpolation argument, combined with some additional considerations specific to the case $n=2$, will yield the estimates for the intermediate values of $p$ in the region where $|\mathrm{Im}\,\sqrt{z}|\geq 1$ (see also [12, Remark 9]). Note that the estimates are uniform in this region, as $\sigma(q)-\frac{1}{2}\leq 0$. We will see that these uniform resolvent estimates account for the second set on the right-hand side of (3). In contrast, the first set, consisting of a union of disks, corresponds to non-uniform resolvent estimates in the complementary region:
(17) $\|(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\lesssim\mathrm{dist}(z,\operatorname{spec}(-\Delta_{g}))^{-1}(1+|z|)^{\sigma(q)},\quad$ $|\mathrm{Im}\,\sqrt{z}|<1.$
These estimates are new in this generality. They are proved by using Sogge’s spectral cluster bounds [20], in the spirit of Bourgain, Shao, Sogge, and Yao [4] and the author’s own work [7]. The “universal” bounds of Sogge [20], valid for any closed manifold $M$, state that
(18) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\lambda-1/2,\lambda+1/2])\|_{L^{2}(M)\to L^{p^{\prime}}(M)}\lesssim(1+\lambda)^{\sigma(q)}.$
Sogge’s bounds hold for the larger range $2\leq p^{\prime}\leq\infty$ (see (20), (34)) than indicated by (4), (15). However, when using (18) to prove resolvent estimates, the latter restrictions are imposed by Sobolev embedding.
We remark that, in the special case $M=S^{n}$, $n\geq 3$ and $p=2n/(n+2)$, (17) was proved in [4, Theorem 1.1]. In fact, they showed that (17) is an approximate equality in this case. This shows that the region of $z\in{\mathbb{C}}$ for which (16) holds for $n\geq 3$ and $p=2n/(n+2)$, is essentially optimal on the sphere and, more generally, on any Zoll manifold. Similarly, optimality of the region $|\mathrm{Im}\,\sqrt{z}|>1$ in (16) for Zoll manifolds, in the case $p=2(n+1)/(n+3)$ and $n\geq 2$, was proved in [12].
Notation
We write $A\lesssim B$ for two non-negative quantities $A,B\geq 0$ to indicate that there is a constant $C>0$ such that $A\leq CB$. If $C$ depends on a list of parameters $\ell$, we sometimes emphasize this by writing $A\lesssim_{\ell}B$. However, the dependence of $C$ on fixed parameters like the manifold $(M,g)$, the dimension $n$ and the Lebesgue exponent $q$ is usually omitted. The notation $A\approx B$ means $A\lesssim B\lesssim A$. We will often consider the frequency $\lambda_{k}$ as an asymptotic parameter, $\lambda_{k}\to\infty$ (thus also $k\to\infty$). The notation $A\ll B$ means that $A\leq cB$, where $c>0$ is suitably small but fixed (in particular, independent of $k$). Sometimes, this notation is used to shorten a “there exists…” statement, e.g. in Theorem 1.3 $k\gg 1$ means “there exists $k_{0}\in{\mathbb{N}}$ such that for all $k\geq k_{0}$, we have…”. We denote the $L^{p}$ norm of a function $f$ on $M$, with respect to the Riemannian measure ${\rm d}x$, by $\|f\|_{L^{p}(M)}$. If $T:L^{s}(M)\to L^{r}(M)$ is a bounded linear operator, we denote its operator norm by $\|T\|_{L^{s}(M)\to L^{r}(M)}$. If $r=s=2$, we sometimes write $\|T\|:=\|T\|_{L^{2}(M)\to L^{2}(M)}$. For a subset $\Omega\subset{\mathbb{R}}$, we define $\mathbf{1}(\sqrt{-\Delta_{g}}\in\Omega)$ by functional calculus. Explicitly, if we denote the orthogonal projection onto the eigenfunction $e_{j}$ by $E_{j}$, that is $E_{j}f=\langle e_{j},f\rangle e_{j}$, then $\mathbf{1}(\sqrt{-\Delta_{g}}\in\Omega)f:=\sum_{\lambda_{j}\in\Omega}E_{j}f$. The inner product is defined as $\langle u,v\rangle:=\int_{M}\overline{u}v\,{\rm d}x$. The closed disk centered at $r\in{\mathbb{C}}$, with radius $r>0$, is denoted by $D(z,r)\subset{\mathbb{C}}$.
2. Resolvent estimates
In the following, we will use the abbreviation
$d(z):=\mathrm{dist}(z,\operatorname{spec}(-\Delta_{g})).$
Since $\Delta_{g}$ is self-adjoint, we have
(19) $\|(-\Delta_{g}-z)^{-1}\|_{L^{2}(M)\to L^{2}(M)}=d(z)^{-1}.$
Note that $p=2$ corresponds to $q=\infty$ and $\sigma(\infty)=0$. Since $q$ will not play any role in this section (except in the proof of Theorem 1.1), we will use the notation $\nu(p^{\prime}):=\sigma(q)$ when $1/q=1/p-1/p^{\prime}$. Explicitly,
(20) $\nu(p^{\prime})=\begin{cases}\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p^{\prime}})\quad&\mbox{if }2\leq p^{\prime}\leq\frac{2(n+1)}{n-1},\\ n(\frac{1}{2}-\frac{1}{p^{\prime}})-\frac{1}{2}\quad&\mbox{if }\frac{2(n+1)}{n-1}\leq p^{\prime}\leq\infty.\end{cases}$
Proposition 2.1.
Let $\Delta_{g}$ be the Laplace–Beltrami operator on an $n$-dimensional closed Riemannian manifold $(M,g)$. Let $2n/(n+2)\leq p\leq 2$ if $n\geq 3$ and $1<p\leq 2$ if $n=2$. Then there is a constant $C=C(M,g,p)$ such that for all $z\in{\mathbb{C}}\setminus\operatorname{spec}(-\Delta_{g})$,
(21) $\|(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}$ $\leq C|z|^{\nu(p^{\prime})-\frac{1}{2}},\quad|\mathrm{Im}\,\sqrt{z}|\geq 1,$
(22) $\|(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}$ $\leq Cd(z)^{-1}(1+|z|)^{\nu(p^{\prime})},\quad|\mathrm{Im}\,\sqrt{z}|<1.$
Remark 2.2.
(i) The region $|\mathrm{Im}\,\sqrt{z}|\geq 1$ corresponds to the set
(23) $\Xi:=\{z\in{\mathbb{C}}:(\mathrm{Im}\,z)^{2}\geq 4(\mathrm{Re}\,z+1)\},$
which is the exterior of a parabola.
(ii) The new result here is (22), whereas (21) is essentially known.
Proof of (21).
If $p=2n/(n+2)$ and $n\geq 3$, or if $2n/(n+2)<p\leq 2(n+1)/(n+3)$ and $n\geq 2$, then (21) follows from [8], or from [12, Theorem 1 and Remark 9], respectively. For $\mathrm{Re}\,z\leq 0$, we have $d(z)=|z|$, so (19) is stronger than (21) when $p=2$. For $\mathrm{Re}\,z\geq 0$, $\mathrm{Im}\,\sqrt{z}\geq 1$, we have $|z|\lesssim d(z)^{2}$, hence (19) implies (21) in this region as well when $p=2$. The Riesz–Thorin interpolation theorem between $p=2n/(n+2)$, $p=2(n+1)/(n+3)$ and $p=2$ then yields (21) for $2n/(n+2)\leq p\leq 2$ and $n\geq 3$. For $n=2$, one interpolates with some $p>1$ arbitrarily close to $p=1$. ∎
Before proceeding with the proof of Proposition 2.1, we record the following simple lemma.
Lemma 2.3.
For any $A\geq 1$ there exists a constant $C_{A}$ (also depending on $n,q$) such that
(24) $\|(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\leq C_{A}d(z)^{-1},\quad\mathrm{Re}\,z\leq A.$
Proof.
We employ a spectral decomposition into frequencies $\lambda_{j}\leq\sqrt{2A}$ and $\lambda_{j}>\sqrt{2A}$. For the latter, we have a uniform bound
(25) $\|\mathbf{1}(-\Delta_{g}>2A)(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\leq C_{A},\quad\mathrm{Re}\,z\leq A.$
Indeed, since for $\mathrm{Re}\,z\leq A$,
(26) $\|\mathbf{1}(-\Delta_{g}>2A)(-\Delta_{g}+1)(-\Delta_{g}-z)^{-1}\|_{L^{2}(M)\to L^{2}(M)}\leq\sup_{\tau>2A}\frac{\tau+1}{|\tau-z|}\lesssim 1,$
it follows from Sobolev embedding (and its dual) that if $f$ has frequencies $\lambda_{j}>\sqrt{2A}$ (i.e., $f$ is in the range of $\mathbf{1}(-\Delta_{g}>2A)$), then
(27) $\|f\|_{L^{p^{\prime}}(M)}$ $\lesssim\|(-\Delta_{g}+1)^{\frac{n}{2}(\frac{1}{2}-\frac{1}{p^{\prime}})}f\|_{L^{2}(M)}$
(28) $\lesssim\|(-\Delta_{g}+1)^{\frac{n}{2}(\frac{1}{2}-\frac{1}{p^{\prime}})-1}(-\Delta_{g}-z)f\|_{L^{2}(M)}$
(29) $\lesssim\|(-\Delta_{g}-z)f\|_{L^{p}(M)}.$
Next, if $f$ has frequencies $\lambda_{j}\leq\sqrt{2A}$, then by Young’s inequality, $\|f\|_{L^{r}(M)}\lesssim_{A}\|f\|_{L^{s}(M)}$ for all $1\leq s\leq r\leq\infty$ (see e.g. [4, Lemma 3.1]). In particular, if $(e_{j})$ is an orthonormal basis of eigenfunctions of $\Delta_{g}$ in $L^{2}(M)$ with eigenvalues $\lambda_{j}^{2}$ and $E_{j}$ denotes the orthogonal projection $E_{j}f=\langle e_{j},f\rangle e_{j}$, then
(30) $\|E_{j}f\|_{L^{p^{\prime}}(M)}\leq\|e_{j}\|_{L^{p^{\prime}}(M)}^{2}\|f\|_{L^{p}(M)}\lesssim\|f\|_{L^{p}(M)}$
for $\lambda_{j}\leq\sqrt{2A}$ and $E_{j}f=0$ for $\lambda_{j}>\sqrt{2A}$. Thus,
(31) $\|(-\Delta_{g}-z)^{-1}f\|_{L^{p^{\prime}}(M)}\leq\sum_{\lambda_{j}\leq\sqrt{2A}}\frac{1}{|\lambda_{j}^{2}-z|}\|E_{j}f\|_{L^{p^{\prime}}(M)}\lesssim_{A}d(z)^{-1}\|f\|_{L^{p}(M)}.$
In the last inequality, we also used that $\{j:\lambda_{j}\leq\sqrt{2A}\}\lesssim_{A}1$ since the spectrum is discrete. ∎
We now turn to the proof of (22).
Proof of (22).
If $\mathrm{Re}\,\sqrt{z}\leq 10$, then (22) follows from Lemma 2.3, so it remains to consider the case $\mathrm{Re}\,\sqrt{z}>10$. We recall that we are also assuming $|\mathrm{Im}\,\sqrt{z}|<1$. We will split the resolvent into two pieces, one involving frequencies close to $\mathrm{Re}\,\sqrt{z}$ and the other involving frequencies at least a unit distance from $\mathrm{Re}\,\sqrt{z}$, i.e. $|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|\leq\delta$ for the first piece and $|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|>\delta$ for the second piece. Here, $0<\delta\leq 1$ is fixed, and we will take $\delta=1/2$ throughout (any other choice can be reduced to this one by scaling the metric). We will prove
(32) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\lesssim d(z)^{-1}|z|^{\nu(p^{\prime})},$
(33) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\lesssim|z|^{\nu(p^{\prime})-\frac{1}{2}}.$
Combining these estimates yields the claimed bound (22). Note that (33) is stronger since $|z|^{-1/2}\lesssim d(z)^{-1}$ for $\mathrm{Im}\,\sqrt{z}<1$, $\mathrm{Re}\,\sqrt{z}>10$.
The proof of (32) uses Sogge’s spectral cluster bounds [20],
(34) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\lambda-1/2,\lambda+1/2])\|_{L^{2}(M)\to L^{p^{\prime}}(M)}\lesssim(1+\lambda)^{\nu(p^{\prime})},$
(35) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\lambda-1/2,\lambda+1/2])\|_{L^{p}(M)\to L^{2}(M)}\lesssim(1+\lambda)^{\nu(p^{\prime})},$
with $\lambda=\mathrm{Re}\,\sqrt{z}$. Thus, by (34), (35), (19),
(36) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}f\|_{L^{p^{\prime}}(M)}$
(37) $\lesssim(\mathrm{Re}\,\sqrt{z})^{\nu(p^{\prime})}\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}f\|_{L^{2}(M)}$
(38) $\leq d(z)^{-1}(\mathrm{Re}\,\sqrt{z})^{\nu(p^{\prime})}\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])f\|_{L^{2}(M)}$
(39) $\lesssim d(z)^{-1}(\mathrm{Re}\,\sqrt{z})^{2\nu(p^{\prime})}\|f\|_{L^{p}(M)},$
which proves (32) since $\mathrm{Re}\,\sqrt{z}\leq|z|^{1/2}$. A more general version of the above argument can be found in [4, Lemma 2.3] and [7, Lemma 3.1].
The proof of (33) is reduced to the case $|\mathrm{Im}\,\sqrt{z}|\geq 1$ by the following argument (similar to the proof of [4, (2.33)]: Choose $z_{1}\in{\mathbb{C}}$ with $\mathrm{Re}\,\sqrt{z_{1}}=\mathrm{Re}\,\sqrt{z}$ and $|\mathrm{Im}\,\sqrt{z_{1}}|=1$. Then
(40) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}f\|_{L^{p^{\prime}}(M)}$
(41) $\leq\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z_{1})^{-1}f\|_{L^{p^{\prime}}(M)}$
(42) $+\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])[(-\Delta_{g}-z)^{-1}-(-\Delta_{g}-z_{1})^{-1}]f\|_{L^{p^{\prime}}(M)}=I+II.$
For the first term $I$, we have
(43) $I$ $\leq\|(-\Delta_{g}-z_{1})^{-1}f\|_{L^{p^{\prime}}(M)}+\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z_{1})^{-1}f\|_{L^{p^{\prime}}(M)}$
(44) $=\mathcal{I}+\mathcal{II}.$
Since $\mathrm{Im}\,\sqrt{z_{1}}=1$, it follows from (21) that $\mathcal{I}$ is bounded by the right-hand side of (33). For $II$ and $\mathcal{II}$, we will again use Sogge’s bounds (34), (35) to show that
(45) $II+\mathcal{II}\lesssim|z|^{\nu(p^{\prime})-\frac{1}{2}}\|f\|_{L^{p}(M)}.$
The proof of the bound for $\mathcal{II}$ is similar to that of (32). The only difference is that the $L^{2}$ bound in the third line of (36) is replaced by
(46) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z_{1})^{-1}f\|_{L^{2}(M)}\lesssim|z|^{-\frac{1}{2}}\|f\|_{L^{2}(M)},$
where we used that, for $\mathrm{Im}\,\sqrt{z_{1}}=1$, $\mathrm{Re}\,\sqrt{z_{1}}>10$,
(47) $|\tau^{2}-z_{1}|=|\tau-\sqrt{z_{1}}||\tau+\sqrt{z_{1}}|\geq(\mathrm{Im}\,\sqrt{z_{1}})(\mathrm{Re}\,\sqrt{z_{1}})\geq\mathrm{Re}\,\sqrt{z}\gtrsim|z|^{\frac{1}{2}}.$
Given the resolvent identity
(48) $(-\Delta_{g}-z)^{-1}-(-\Delta_{g}-z_{1})^{-1}=(z-z_{1})(-\Delta_{g}-z)^{-1}(-\Delta_{g}-z_{1})^{-1}$
and the fact that $|z-z_{1}|\lesssim\mathrm{Re}\,\sqrt{z}$, the bound for $II$ in (45) would follow from
(49) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}(-\Delta_{g}-z_{1})^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\lesssim|z|^{\nu(p^{\prime})-1}.$
Since the left-hand side is dominated by
(50) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-z)^{-1}\|_{L^{2}(M)\to L^{p^{\prime}}(M)}\|(-\Delta_{g}-z_{1})^{-1}\|_{L^{p}(M)\to L^{2}(M)},$
and since the $L^{2}(M)\to L^{p^{\prime}}(M)$ norm equals the $L^{p}(M)\to L^{2}(M)$ norm of the adjoint operator, it suffices to prove
(51) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2])(-\Delta_{g}-\overline{z})^{-1}\|_{L^{p}(M)\to L^{2}(M)}$ $\lesssim|z|^{\frac{\nu(p^{\prime})}{2}-\frac{1}{2}},$
(52) $\|(-\Delta_{g}-z_{1})^{-1}\|_{L^{p}(M)\to L^{2}(M)}$ $\lesssim|z|^{\frac{\nu(p^{\prime})}{2}-\frac{1}{2}}.$
We start with the proof of (52). Since $\mathrm{Im}\,\sqrt{z_{1}}=1$ and $\mathrm{Re}\,\sqrt{z_{1}}=\mathrm{Re}\,\sqrt{z}$, we have the lower bound
(53) $|\lambda_{j}^{2}-z_{1}|=|\lambda_{j}+\sqrt{z_{1}}||\lambda_{j}-\sqrt{z_{1}}|\gtrsim(\mathrm{Re}\,\sqrt{z})(|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|+1),$
Thus, by $L^{2}$ orthogonality,
(54) $\|(-\Delta_{g}-z_{1})^{-1}f\|_{L^{2}(M)}^{2}$ $=\sum_{j=0}^{\infty}|\lambda_{j}^{2}-z_{1}|^{-2}\|E_{j}f\|_{L^{2}(M)}^{2}$
(55) $\lesssim(\mathrm{Re}\,\sqrt{z})^{-2}\sum_{j=0}^{\infty}(|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|+1)^{-2}\|E_{j}f\|_{L^{2}(M)}^{2}$
(56) $\lesssim(\mathrm{Re}\,\sqrt{z})^{-2}\sum_{k=1}^{\infty}\sup_{\tau\in[k-1,k)}(|\tau-\mathrm{Re}\,\sqrt{z}|+1)^{-2}\|\sum_{\lambda_{j}\in[k-1,k)}E_{j}f\|_{L^{2}(M)}^{2}$
(57) $\lesssim(\mathrm{Re}\,\sqrt{z})^{-2}\|f\|_{L^{p}(M)}^{2}\sum_{k=1}^{\infty}(|k-\mathrm{Re}\,\sqrt{z}|+1)^{-2}k^{2\nu(p^{\prime})}.$
In the last line, we used Sogge’s spectral cluster bound (35). We split the sum into two pieces, one with $|k-\mathrm{Re}\,\sqrt{z}|\leq\frac{1}{2}\mathrm{Re}\,\sqrt{z}$ and the other with $|k-\mathrm{Re}\,\sqrt{z}|>\frac{1}{2}\mathrm{Re}\,\sqrt{z}$. For the first piece, we use $k\leq\frac{3}{2}\mathrm{Re}\,\sqrt{z}$ to estimate
(58) $\sum_{|k-\mathrm{Re}\,\sqrt{z}|\leq\frac{1}{2}\mathrm{Re}\,\sqrt{z}}(|k-\mathrm{Re}\,\sqrt{z}|+1)^{-2}k^{2\nu(p^{\prime})}\lesssim(\mathrm{Re}\,\sqrt{z})^{2\nu(p^{\prime})}.$
For the second piece, we use that $|k-\mathrm{Re}\,\sqrt{z}|+1>\frac{1}{2}|k-\mathrm{Re}\,\sqrt{z}|+\frac{1}{4}\mathrm{Re}\,\sqrt{z}$, leading to the better estimate
(59) $\sum_{|k-\mathrm{Re}\,\sqrt{z}|>\frac{1}{2}\mathrm{Re}\,\sqrt{z}}(|k-\mathrm{Re}\,\sqrt{z}|+1)^{-2}k^{2\nu(p^{\prime})}$
(60) $\lesssim\int_{-\infty}^{\infty}(|k-\mathrm{Re}\,\sqrt{z}|+\mathrm{Re}\,\sqrt{z})^{-2}k^{2\nu(p^{\prime})}{\rm d}k\lesssim(\mathrm{Re}\,\sqrt{z})^{2\nu(p^{\prime})-1}.$
Altogether, we obtain (52).
For the proof of (51), we use that $\mathrm{Re}\,\sqrt{\overline{z}}=\mathrm{Re}\,\sqrt{z}$ and $\mathrm{Im}\,\sqrt{\overline{z}}=-\mathrm{Im}\,\sqrt{z}$. Moreover, for $\lambda_{j}\notin[\mathrm{Re}\,\sqrt{z}-1/2,\mathrm{Re}\,\sqrt{z}+1/2]$,
(61) $|\lambda_{j}^{2}-\overline{z}|=|\lambda_{j}+\sqrt{\overline{z}}||\lambda_{j}-\sqrt{\overline{z}}|\gtrsim(\mathrm{Re}\,\sqrt{z})(|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|+1).$
Thus, (51) follows by repeating the proof of (52). ∎
We are now ready to prove Theorem 1.1. For the proof, we will revert to the notation $\sigma(q)$, recalling that $\nu(p^{\prime})=\sigma(q)$ when $1/q=1/p-1/p^{\prime}$.
Proof of Theorem 1.1.
By the Birman–Schwinger principle, $z\in{\mathbb{C}}\setminus\operatorname{spec}(-\Delta_{g})$ is an eigenvalue of $\Delta_{g}+V$ if and only if 1-1 is an eigenvalue of the (compact) Birman–Schwinger operator $K(z):=|V|^{1/2}(-\Delta_{g}-z)^{-1}V^{1/2}$. In particular, if $z$ is an eigenvalue, then $K(z)$ has norm at least one, which implies that
(62) $1\leq\|K(z)\|\leq\|V\|_{L^{q}(M)}\|(-\Delta_{g}-z)^{-1}\|_{L^{p}(M)\to L^{p^{\prime}}(M)},$
where $1/q=1/p-1/p^{\prime}$. Thus, by (62) and the resolvent estimates (21), (22), if $z\in\operatorname{spec}(-\Delta_{g}+V)$, then either
(63) $|z|^{\frac{1}{2}}(1+|z|)^{-\sigma(q)}\lesssim\|V\|_{L^{q}(M)}\quad\mbox{or}\quad d(z)\lesssim(1+|z|)^{\sigma(q)}\|V\|_{L^{q}(M)}.$
In the first case, (3) trivially follows. The second case in (63) occurs for $|\mathrm{Im}\,\sqrt{z}|<1$, or equivalently, $z\in\Xi^{c}=\{z\in{\mathbb{C}}:(\mathrm{Im}\,z)^{2}<4(\mathrm{Re}\,z+1)\}$, see (23). The set $\Xi^{c}$ is contained in the union of a neighborhood of the origin and a parabolic region around $(1,\infty)$,
(64) $\Xi^{c}\subset\{z\in{\mathbb{C}}:|z|<\sqrt{8}\}\cup\{z\in{\mathbb{C}}:\mathrm{Re}\,z>1,\,(\mathrm{Im}\,z)^{2}<8\mathrm{Re}\,z\}.$
If $|z|<\sqrt{8}$, then the second inequality in (63) implies that $z\in D(\lambda_{0}^{2},Cr_{0})$ (recalling that $\lambda_{0}=0$). If $\mathrm{Re}\,z>1,\,(\mathrm{Im}\,z)^{2}<8\mathrm{Re}\,z$, then there exists a unique $k\in{\mathbb{N}}$ such that $\mathrm{Re}\,z\in[\lambda_{k-1}^{2},\lambda_{k}^{2})$. The second inequality in (63) then implies that $d(z)\leq Cr_{k}$. Since $d(z)=\min\{|z-\lambda_{k-1}^{2}|,|z-\lambda_{k}^{2}|\}$ and since $\lim_{j\to\infty}\frac{\lambda_{j}}{\lambda_{j-1}}=1$ by Weyl’s asymptotics, it follows that $z\in D(\lambda_{k-1}^{2},r_{k-1})\cup D(\lambda_{k}^{2},r_{k})$. This completes the proof of (3), and hence of Theorem 1.1 in the case $q>n/2$.
If $n\geq 3$ and $q=n/2$, then (21) yields a uniform bound, i.e.
(65) $\|(-\Delta_{g}-z)^{-1}\|_{L^{\frac{2n}{n+2}}(M)\to L^{\frac{2n}{n-2}}(M)}$ $\leq C_{0},\quad|\mathrm{Im}\,\sqrt{z}|\geq 1.$
Therefore, (62) implies that $\|V\|_{L^{\frac{n}{2}}(M)}\geq C_{0}^{-1}$. Put differently, $z$ cannot be an eigenvalue if $|\mathrm{Im}\,\sqrt{z}|\geq 1$ and $\|V\|_{L^{\frac{n}{2}}(M)}<C_{0}^{-1}$. Thus, if we set $c:=C_{0}^{-1}/2$, then all eigenvalues of $\Delta_{g}+V$, $\|V\|_{L^{\frac{n}{2}}(M)}\leq c$, are contained in the region $|\mathrm{Im}\,\sqrt{z}|<1$. Using (22), we obtain the refinement stated in (5). ∎
3. Optimality of eigenvalue bounds
In this section, we prove Theorem 1.3. We will seek the potential in the form
(66) $V=\kappa\chi^{2},\quad 0\leq\chi=\chi_{k}\in L^{2q}(S^{n}),\quad\|\chi\|_{L^{2q}(S^{n})}=1,\quad\kappa=\kappa_{k,\theta,\rho}\in{\mathbb{C}},\quad 0\leq|\kappa|\lesssim\rho.$
3.1. Optimality on $S^{n}$
We recall that
(67) $\operatorname{spec}(-\Delta_{S^{n}})=\{\lambda_{k}^{2}=k(k+n-1):k=0,1,2,\ldots\}$
and that the eigenspace $H_{k}$ corresponding to the eigenvalue k(k+n1)k(k+n-1), the space of spherical harmonics of degree $k$, has dimension
(68) $d_{k}:=\dim H_{k}=\begin{pmatrix}n+k\\ n\end{pmatrix}-\begin{pmatrix}n+k-2\\ n\end{pmatrix}\approx k^{n-1},$
see e.g. [18, Section 22.3] or [21, Section 3.4].
Step 1: By the Birman–Schwinger principle (see the first paragraph in the proof of Theorem 1.1), it suffices to find $\chi,\kappa$ as above such that $\kappa^{-1}$ is an eigenvalue of the compact operator $\chi(-\Delta_{S^{n}}-z_{k,\theta,\rho})^{-1}\chi$, where
(69) $z_{k,\theta,\rho}=k(k+n-1)+\rho{\rm e}^{\mathrm{i}\theta}(1+\mathcal{O}(\rho k^{2\sigma(q)-1}))k^{2\sigma(q)}.$
In other words, we seek a zero $(z,\kappa)=(z_{k,\theta,\rho},\kappa_{k,\theta,\rho})$ of the operator-valued function
(70) $\mathcal{A}(z,\kappa):=I+\kappa\chi(-\Delta_{S^{n}}-z)^{-1}\chi,\quad z\in{\mathbb{C}}\setminus\operatorname{spec}(-\Delta_{S^{n}}),\quad|\kappa|\lesssim\rho.$
For fixed $\kappa$, we denote
(71) $\operatorname{spec}{\mathcal{A}(\cdot,\kappa)}:=\{z\in{\mathbb{C}}:\mathcal{A}(z,\kappa)\mbox{ is not invertible in }L^{2}(S^{n})\}.$
For $z\in D_{k}$ (defined in (10)), we decompose $\chi(-\Delta_{S^{n}}-z)^{-1}\chi$ into a ‘resonant’ part and a ‘nonresonant’ part,
(72) $K_{\rm r}(z):=\frac{\chi P_{k}\chi}{k(k+n-1)-z},\quad K_{\rm nr}(z):=\chi(-\Delta_{S^{n}}-z)^{-1}\chi-K_{\rm r}(z),$
where $P_{k}$ is the orthogonal projection onto $H_{k}$. Note that
(73) $|\sqrt{z}-\sqrt{k(k+n-1)}|\lesssim\rho k^{2\sigma(q)-1}\quad\mbox{for all }z\in D_{k}.$
Thus, if $q>n/2$ and $k\gg 1+\rho^{1/(1-2\sigma(q))}$, or if $q=n/2$ and $\rho$ is sufficiently small, we have
(74) $|\sqrt{z}-\sqrt{k(k+n-1)}|<1/2,\quad\min_{j\neq k}|\sqrt{z}-\sqrt{j(j+n-1)}|\geq 1/2,$
for all $z\in D_{k}$; the second inequality also uses that $|\sqrt{k(k+n-1)}-\sqrt{j(j+n-1)}|\gtrsim 1$ for $j\neq k$. Then (73) implies that we may write
(75) $K_{\rm r}(z)$ $=\mathbf{1}(\sqrt{-\Delta_{S^{n}}}\in[\sqrt{k(k+n-1)}-1/2,\sqrt{k(k+n-1)}+1/2])(-\Delta_{S^{n}}-z)^{-1},$
(76) $K_{\rm nr}(z)$ $=\mathbf{1}(\sqrt{-\Delta_{S^{n}}}\notin[\sqrt{k(k+n-1)}-1/2,\sqrt{k(k+n-1)}+1/2])(-\Delta_{S^{n}}-z)^{-1},$
and hence, by (32), (33), together with Hölder’s inequality and the normalization condition on $\chi$ in (66), we have the bounds
(77) $\sup_{z\in D_{k}}\|K_{\rm r}(z)\|_{L^{2}(S^{n})\to L^{2}(S^{n})}$ $\lesssim k^{2\sigma(q)}|k(k+n-1)-z|^{-1},$
(78) $\sup_{z\in D_{k}}\|K_{\rm nr}(z)\|_{L^{2}(S^{n})\to L^{2}(S^{n})}$ $\lesssim k^{2\sigma(q)-1}.$
We will choose $\chi$ so that (77) becomes an approximate equality.
Step 2: Since $P_{k}$ has rank $d_{k}$ and $\chi$ is bounded, $\chi P_{k}\chi$ has rank at most $d_{k}$; in particular, it is a compact operator. Moreover, for any $f\in L^{2}(S^{n})$, $\langle\chi P_{k}\chi f,f\rangle=\|P_{k}\chi f\|^{2}_{L^{2}(S^{n})}\geq 0$, thus $\chi P_{k}\chi$ is self-adjoint and non-negative, having at most $d_{k}$ nonzero eigenvalues $a_{j}(k)$ (counted with multiplicity),
(79) $\|\chi P_{k}\chi\|=a_{0}(k)\geq a_{1}(k)\geq\ldots\geq a_{d_{k}}(k)\geq 0.$
By the spectral theorem for compact, self-adjoint operators, $\chi P_{k}\chi$ has an orthonormal basis of eigenfunctions $\{f_{k,i}\}_{i=1}^{d_{k}}$. If $\{e_{k,i}\}_{i=1}^{d_{k}}$ is an orthonormal basis of $\mathop{\rm Ran}P_{k}$ (thus, $e_{k,i}$ are eigenfuncitons of $\Delta_{g}$ corresponding to eigenvalue $\lambda_{k}^{2}$), then the functions $f_{k,j}$ lie in the linear span of $\{\chi e_{k,i}\}_{i=1}^{d_{k}}$. Moreover, the eigenvalues $a_{j}(k)$ can be computed by diagonalizing the Gram matrix $G=(G_{ij})$, with $G_{ij}=\langle\chi e_{k,i},\chi e_{k,j}\rangle$, $1\leq i,j\leq d_{k}$.
Lemma 3.1.
Let $n\geq 2$, $q>n/2$ or $n\geq 3$, $q\geq n/2$. For $k\in{\mathbb{N}}$ sufficiently large, there exists a nonnegative function $\chi=\chi_{k}\in L^{2q}(S^{n})$, $\|\chi\|_{L^{2q}(S^{n})}=1$, and a constant $C_{a_{0}}>0$, such that
(80) $C_{a_{0}}^{-1}k^{2\sigma(q)}\leq a_{0}(k)\leq C_{a_{0}}k^{2\sigma(q)}.$
Proof.
Sogge proved [19, Theorem 4.1] that
(81) $\|P_{k}\|_{L^{p}(S^{n})\to L^{2}(S^{n})}\approx k^{\nu(p^{\prime})},$
where we recall that $\nu(p^{\prime})=\sigma(q)$ when $1/q=1/p-1/p^{\prime}$. By a $TT^{*}$ argument, we also have
(82) $\|P_{k}\|_{L^{p}(S^{n})\to L^{p^{\prime}}(S^{n})}\approx k^{2\nu(p^{\prime})}.$
The upper bound in (82) and Hölder’s inequality imply that for any nonnegative function $\chi$ with $\|\chi\|_{L^{2q}(S^{n})}=1$,
(83) $\|\chi P_{k}\chi\|=\sup_{\|f\|_{L^{2}(S^{n})}=1}\langle f,\chi P_{k}\chi f\rangle\leq\|P_{k}\|_{L^{p}(S^{n})\to L^{p^{\prime}}(S^{n})}\sup_{\|f\|_{L^{2}(S^{n})}=1}\|\chi f\|_{L^{p}(S^{n})}\lesssim k^{2\sigma(q)}$
since $1/p=1/(2q)+1/2$.
If we take the supremum of (83) over all normalized $\chi\in L^{2q}(S^{n})$, we obtain
(84) $\sup_{\|\chi\|_{L^{2q}(S^{n})}=1}\|\chi P_{k}\chi\|=\sup_{\begin{subarray}{c}\|\chi\|_{L^{2q}(S^{n})}=1\\ \|f\|_{L^{2}(S^{n})}=1\end{subarray}}\langle\chi f,P_{k}\chi f\rangle\geq\sup_{\|g\|_{L^{p}(S^{n})}=1}\langle g,P_{k}g\rangle=\|P_{k}\|_{L^{p}(S^{n})\to L^{p^{\prime}}(S^{n})},$
where the inequality follows from the fact that any normalized $g\in L^{p}(S^{n})$ can be factored as $g=\chi f$, where $\chi$ and $f$ are normalized functions in $L^{2q}(S^{n})$ and $L^{2}(S^{n})$, respectively. Indeed, if we write $g={\rm e}^{\mathrm{i}\arg g}|g|$, then we can take $\chi=|g|^{\frac{p}{2q}}$ and $f={\rm e}^{\mathrm{i}\arg g}|g|^{\frac{p}{2}}$. It follows from (82) that there exists a normalized function $\chi\in L^{2q}(S^{n})$ such that we have $\|\chi P_{k}\chi\|\gtrsim k^{2\sigma(q)}$. The lemma is proved since the supremum can be taken over nonnegative functions $\chi$. ∎
Remark 3.2.
(i) The function $\chi$ can be constructed as follows. By (81), there exists $\psi_{k}\in L^{p}(S^{n})$ such that
(85) $\frac{\|P_{k}\psi_{k}\|_{L^{2}(S^{n})}}{\|\psi_{k}\|_{L^{p}(S^{n})}}\approx k^{\sigma(q)},$
and we normalize $\psi_{k}$ so that $\|\psi_{k}\|_{L^{p}(S^{n})}=1$. With this choice, we have
(86) $\langle\psi_{k},P_{k}\psi_{k}\rangle=\|P_{k}\psi_{k}\|_{L^{2}(S^{n})}^{2}\approx k^{2\sigma(q)}.$
Similarly as in the proof of Lemma 3.1, setting $\chi_{k}=|\psi_{k}|^{\frac{p}{2q}}$ and $f_{k}={\rm e}^{\mathrm{i}\arg\psi_{k}}|\psi_{k}|^{\frac{p}{2}}$, we then have $\psi_{k}=\chi_{k}f_{k}$, $\|\chi_{k}\|_{L^{2q}(S^{n})}=1$, $\|f_{k}\|_{L^{2}(S^{n})}=1$, and
(87) $\|\chi_{k}P_{k}\chi_{k}\|\geq\langle\chi_{k}f_{k},P_{k}\chi_{k}f_{k}\rangle\approx k^{2\sigma(q)}.$
We refer to [22, Chapter 5.1] for the construction of $\psi_{k}$ in (85), which depends on whether $2(n+1)/(n-1)\leq p^{\prime}\leq\infty$ or $2\leq p^{\prime}\leq 2(n+1)/(n-1)$. This construction works on any compact manifold and shows that Sogge’s unit-width spectral cluster bounds (34), (35) are always optimal.
(ii) On $S^{n}$, the construction is even more explicit, at least when $2\leq p^{\prime}\leq 2(n+1)/(n-1)$. Sogge [19] showed that in this case, (85) is saturated by the highest weight spherical harmonic $Q_{k}$ of degree $k$. Then $\chi_{k}=|Q_{k}|^{\frac{p}{2q}}$ is completely explicit. Sogge also showed that the dual of (81) is saturated by the $k$-th zonal spherical harmonic $Z_{k}$ for $2(n+1)/(n-1)\leq p^{\prime}\leq\infty$, but $Z_{k}$ cannot be directly used in (85) since it does not have the stated $L^{2}$ to $L^{p}$ norm ratio in that range of $p$.
Step 3: We will compare the zeros of $\mathcal{A}(\cdot,\kappa)$ (defined in (70)) to the zeros $z_{{\rm r},j}(\kappa)$ of the following operator-valued function,
(88) $\mathcal{A}_{\rm r}(z,\kappa):=I+\kappa K_{\rm r}(z),\quad z\in{\mathbb{C}}\setminus\operatorname{spec}(-\Delta_{S^{n}}),\quad|\kappa|\lesssim\rho,$
where $K_{\rm r}$ was defined in (72). These zeros are given by
(89) $\operatorname{spec}(\mathcal{A}_{\rm r}(z,\kappa))=\{z_{{\rm r},j}(\kappa):=k(k+n-1)+\kappa a_{j}(k):j=0,\ldots,d_{k}\}.$
Setting
(90) $\kappa_{0}:=\rho{\rm e}^{\mathrm{i}\theta}k^{2\sigma(q)}a_{0}(k)^{-1}\in D(0,C_{a_{0}}\rho),$
we find that
(91) $z_{{\rm r},0}(\kappa_{0})=k(k+n-1)+\rho{\rm e}^{\mathrm{i}\theta}k^{2\sigma(q)},\quad|z_{{\rm r},0}(\kappa_{0})-z_{k,\theta,\rho}|\lesssim\rho^{2}k^{4\sigma(q)-1},$
where $z_{k,\theta,\rho}$ is given by (69). Hence, by the triangle inequality, the proof of the first part of Theorem 1.3 would be complete if we could show that
(92) $D(z_{{\rm r},0}(\kappa_{0}),C\rho^{2}k^{4\sigma(q)-1})\cap\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\neq\emptyset$
for sufficiently large $k$ and some universal constant $C$.
To this end, we will use an operator version of Rouché’s theorem, due to Gohberg and Sigal [13, Theorem 2.2]. It implies that if $\gamma$ is a simple closed rectifiable contour bounding a domain $U\subset{\mathbb{C}}$ and if
(93) $\max_{w\in\gamma}\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}(\mathcal{A}(w,\kappa_{0})-\mathcal{A}_{\rm r}(w,\kappa_{0}))\|<1,$
then
(94) $\frac{1}{2\pi\mathrm{i}}\operatorname{Tr}\oint_{\gamma}\partial_{w}\mathcal{A}_{\rm r}(w,\kappa_{0})\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}{\rm d}w=\frac{1}{2\pi\mathrm{i}}\operatorname{Tr}\oint_{\gamma}\partial_{w}\mathcal{A}(w,\kappa_{0})\mathcal{A}(w,\kappa_{0})^{-1}{\rm d}w.$
Ideally, we would like to choose the contour $\gamma=\partial U$ to be a circle with center $z_{{\rm r},0}(\kappa_{0})$ and radius $C\rho^{2}k^{4\sigma(q)-1}$. However, due to their $k$-dependence, it is not clear how to avoid zeros of $\mathcal{A}_{\rm r}$ and $\mathcal{A}$. The main problem is that we have no uniform (independent of $k$) lower bound on the gap between the largest eigenvalue $a_{0}(k)$ of $\chi P_{k}\chi$ and the next largest distinct eigenvalue (recall that, in (79), the eigenvalues $a_{j}(k)$ were counted with multiplicity.)
Lemma 3.3.
Let $n\geq 2$, $q>n/2$ or $n\geq 3$, $q\geq n/2$, let $\kappa_{0}$, $z_{{\rm r},0}(\kappa_{0})$ be as as in (90), (91), and assume that $\|\chi\|_{L^{2q}(S^{n})}=1$. For $\epsilon,\rho>0$, $\epsilon$ sufficiently small, and $k\in{\mathbb{N}}$ sufficiently large, such that $\rho\epsilon^{-1}k^{2\sigma(q)-1}\ll 1$, there exists a circle $\gamma=\partial U$ with center $z_{{\rm r},0}(\kappa_{0})$ and radius $\mathcal{O}(\epsilon\rho)k^{2\sigma(q)}$, which does not intersect any zeros of $\mathcal{A}_{\rm r}(\cdot,\kappa_{0})$ or $\mathcal{A}(\cdot,\kappa_{0})$, and such that
(95) $\#(\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\cap U)=\#(\operatorname{spec}{\mathcal{A}_{\rm r}(\cdot,\kappa_{0})}\cap U).$
Conclusion of the proof of Theorem 1.3 i).
Since $z_{{\rm r},0}(\kappa_{0})\in\operatorname{spec}{\mathcal{A}_{\rm r}(\cdot,\kappa_{0})}$, the right-hand side of (95) is $\geq 1$. Since we may take $\epsilon=C\rho k^{2\sigma(q)-1}$ for some $C\gg 1$ in Lemma 3.3, it follows that
(96) $\#(\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\cap D(z_{{\rm r},0}(\kappa_{0}),C\rho^{2}k^{4\sigma(q)-1}))\geq 1,$
which means that (92) holds. ∎
Proof of Lemma 3.3.
1. We will use the following bound on the eigenvalues $a_{j}(k)$ of $\chi P_{k}\chi$,
(97) $\Big(\sum_{j}a_{j}(k)^{\beta(q)}\Big)^{1/\beta(q)}\leq C_{\beta(q)}k^{2\sigma(q)},$
for some $\beta(q)\in[1,\infty)$, which is a consequence of the Schatten norm bounds of Frank and Sabin [11]. Specifically, this follows from [11, Remark 6] in the case $M=S^{n}$ and $\lambda\approx k$. The Schatten exponent $\beta(q)$ is given explicitly by
(98) $\beta(q)=\begin{cases}\frac{(n-1)q}{n-q}\quad&\mbox{if }\frac{n}{2}\leq q\leq\frac{n+1}{2},\\ 2q\quad&\mbox{if }\frac{n+1}{2}\leq q\leq\infty,\end{cases}$
however, the exact value of $\beta(q)$ is not important to us here. From (97) and the monotonicity of $a_{j}(k)$, we get
(99) $a_{j}(k)\leq C_{\beta(q)}j^{-1/\beta(q)}k^{2\sigma(q)}\leq C_{\beta(q)}C_{a_{0}}j^{-1/\beta(q)}a_{0}(k),$
where the second estimate follows from Lemma 3.1. Let $j_{0}=j_{0}(q)$ be the first integer such that $j_{0}>(C_{\beta(q)}C_{a_{0}})^{\beta(q)}$. Then
(100) $a_{0}(k)-a_{j_{0}}(k)\geq C_{1}a_{0}(k),\quad\mbox{where}\quad C_{1}:=(1-C_{\beta(q)}C_{a_{0}}j_{0}^{-1/\beta(q)})>0.$
Moreover, there are at most $j_{0}$ eigenvalues $a_{1}(k),\ldots,a_{j_{0}}(k)$ in the interval $a_{j_{0}}(k),a_{0}(k)]$. If we rescale $\tilde{a}_{j}(k):=a_{j}(k)/a_{0}(k)$, then
(101) $0\leq\tilde{a}_{j_{0}}(k)\leq\ldots\leq\tilde{a}_{1}(k)\leq\tilde{a}_{0}(k)=1,\quad\tilde{a}_{j_{0}}(k)\leq 1-C_{1}.$
Let $0<\epsilon<C_{1}/2$, so that $I_{\epsilon}:=[1-2\epsilon,1]$ is a subinterval of $\tilde{a}_{j_{0}}(k),1]$ and thus contains at most $j_{0}$ distinct points of the sequence $\{\tilde{a}_{j}(k)\}_{j=0}^{d_{k}}$. Let
(102) $I_{\epsilon,m}:=[1-\frac{2\epsilon m}{j_{0}+1},1-\frac{2\epsilon(m-1)}{j_{0}+1}],\quad m=1,2,\ldots,j_{0}+1.$
By the pigeonhole principle, there exists at least one interval $I_{\epsilon,m(k)}$ that contains no point of $\{\tilde{a}_{j}(k)\}_{j=0}^{d_{k}}$. Let $\tilde{x}(k)$ be the midpoint of this interval. By construction,
(103) $\min_{0\leq j\leq d_{k}}|\tilde{x}(k)-\tilde{a}_{j}(k)|\geq\frac{\epsilon}{j_{0}+1},\quad 2\epsilon\geq 1-\tilde{x}(k)\geq\frac{\epsilon}{j_{0}+1}.$
We now define
(104) $U:=D(z_{{\rm r},0}(\kappa_{0}),|\kappa_{0}|(1-\tilde{x}(k))a_{0}(k))\subset D_{k},\quad\gamma:=\partial U.$
By (103), (90), we can write
(105) $U=D(z_{{\rm r},0}(\kappa_{0}),\mathcal{O}(\epsilon\rho)k^{2\sigma(q)}).$
2. To prove $\gamma\cap\operatorname{spec}(\mathcal{A}_{\rm r}(\cdot,\kappa_{0}))=\emptyset$, we first recall that
(106) $\operatorname{spec}(\mathcal{A}_{\rm r}(\cdot,\kappa_{0}))=\{z_{{\rm r},j}(\kappa_{0})\}_{j=0}^{d_{k}},$
where $z_{{\rm r},j}(\kappa)$ are given by (89). We fix $j\in\{0,\ldots,d_{k}\}$ and $w\in\gamma$, and write
(107) $w=z_{{\rm r},0}(\kappa_{0})+{\rm e}^{\mathrm{i}\varphi}\kappa_{0}(1-\tilde{x}(k))a_{0}(k))$
for some $\varphi\in[0,2\pi)$. Then
(108) $|w-z_{{\rm r},j}(\kappa_{0})|$ $\geq|w-z_{{\rm r},0}(\kappa_{0})|-|z_{{\rm r},j}(\kappa_{0})-z_{{\rm r},0}(\kappa_{0})|$
(109) $\geq|\kappa_{0}|(1-\tilde{x}(k))a_{0}(k)-|\kappa_{0}|(a_{0}(k)-a_{j}(k))$
(110) $=|\kappa_{0}|(\tilde{a}_{j}(k)-\tilde{x}(k))a_{0}(k)\geq\frac{\epsilon\rho k^{2\sigma(q)}}{j_{0}+1},$
where we used the first bound of (103) and the definition of $\kappa_{0}$, see (90), in the last inequality. This implies that
(111) $\mathrm{dist}(\gamma,\operatorname{spec}(\mathcal{A}_{\rm r}(\cdot,\kappa_{0})))$ $\geq\frac{\epsilon\rho k^{2\sigma(q)}}{j_{0}+1}.$
and thus $\gamma\cap\operatorname{spec}(\mathcal{A}_{\rm r}(\cdot,\kappa_{0}))=\emptyset$.
3. To prove $\gamma\cap\operatorname{spec}(\mathcal{A}(\cdot,\kappa_{0}))=\emptyset$, let $w\in\gamma$, as in (107). Since $K_{\rm r}(w)$ is a normal operator, we have
(112) $\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}\|$ $=\max_{0\leq j\leq d_{k}}|\Big(1+\frac{\kappa_{0}a_{j}(k)}{k(k+n-1)-w}\Big)^{-1}|$
(113) $=|k(k+n-1)-w|\max_{0\leq j\leq d_{k}}|w-z_{{\rm r},j}(\kappa_{0})|^{-1}.$
By (89), (105) and the second formula in (103), we have
(114) $|k(k+n-1)-w|$ $\leq|k(k+n-1)-z_{{\rm r},0}(\kappa_{0})|+|w-z_{{\rm r},0}(\kappa_{0})|$
(115) $\leq|\kappa_{0}|a_{0}(k)+|\kappa_{0}|a_{0}(k)(1-\tilde{x}(k))$
(116) $\leq\rho k^{2\sigma(q)}(1+2\epsilon).$
Together with (108), (112), we then obtain
(117) $\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}\|\leq(1+2\epsilon)(j_{0}+1)\epsilon^{-1}.$
Then, since
(118) $\mathcal{A}(w,\kappa_{0})=\mathcal{A}_{\rm r}(w,\kappa_{0})(I+\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}(\mathcal{A}(w,\kappa_{0}))-\mathcal{A}_{\rm r}(w,\kappa_{0})),$
and, by (78),
(119) $\|\mathcal{A}(w,\kappa_{0})-\mathcal{A}_{\rm r}(w,\kappa_{0})\|=|\kappa_{0}|\|K_{\rm nr}\|\lesssim\rho k^{2\sigma(q)-1},$
it follows that $\mathcal{A}(w,\kappa_{0})$ is invertible, and
(120) $\|\mathcal{A}(w,\kappa_{0})^{-1}\|\leq(1+2\epsilon)(j_{0}+1)\epsilon^{-1}(1-\mathcal{O}(\rho\epsilon^{-1}k^{2\sigma(q)-1}))^{-1}.$
This proves that $\gamma\cap\operatorname{spec}(\mathcal{A}(\cdot,\kappa_{0}))=\emptyset$.
4. Finally, we check that the smallness condition (93) holds. By (117) and (78), we have
(121) $\max_{w\in\gamma}\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}(\mathcal{A}(w,\kappa_{0})-\mathcal{A}_{\rm r}(w,\kappa_{0}))\|$
(122) $\leq|\kappa_{0}|\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}\|\max_{w\in\gamma}\|K_{\rm nr}\|$
(123) $\lesssim\rho\epsilon^{-1}k^{2\sigma(q)-1}\ll 1.$
Now, [13, Theorem 2.2] yields (94), which is equivalent to (95). ∎
3.2. Optimality on Zoll manifolds
The proof of the second part of Theorem 1.3 is analogous to that of the first part, with a few modifications, which we shall point out below. As mentioned in Remark 1.4, optimality for $q=\infty$ is trivial, therefore we exclude this case here.
Step 1: We replace $\Delta_{S^{n}}$ with $\Delta_{g}$ in the definition (70) and $z_{k,\theta,\rho}$ in (69) by
(124) $z_{k,\theta,\rho}=(k+\alpha)^{2}+\rho{\rm e}^{\mathrm{i}\theta}(1+\mathcal{O}(\rho k^{2\sigma(q)-1}))k^{2\sigma(q)}.$
Note that, due to the additional assumption $k^{2\sigma(q)}\rho\gg 1$ in the Zoll case, the eigenvalues $\lambda_{k,j}^{2}$ of $\Delta_{g}$ corresponding to the $k$-th cluster are contained inside the disks $\tilde{D}_{k}$ defined in (12), and these disks are mutually disjoint. The situation is similar to that on the sphere, the only difference being that a single eigenvalue is now replaced by a whole cluster of eigenvalues. For this reason, we use a different splitting $K(z)=K_{\rm r}(z)+K_{\rm nr}(z)$, namely
(125) $K_{\rm r}(z)$ $:=\frac{\chi\mathbf{1}(\sqrt{-\Delta_{g}}\in[k+\alpha-A/k,k+\alpha+A/k])\chi}{(k+\alpha)^{2}-z}$
(126) $=\frac{\chi\mathbf{1}(\sqrt{-\Delta_{g}}\in[k+\alpha-1/2,k+\alpha+1/2])\chi}{(k+\alpha)^{2}-z},$
where the second equality holds due to the clustering of the eigenvalues.
We first establish an analogue of (78). Note that we only used this in the last step of the proof of Lemma 3.3 over the curve $\gamma$, not over the whole disk $D_{k}$.
Lemma 3.4.
Assume that $n/2<q<\infty$ and $k\gg 1$ or $q=n/2$, $n\geq 3$, and $\rho\ll 1$. Then
(127) $\|K_{\rm nr}(z)\|$ $\lesssim|(k+\alpha)^{2}-z|^{-2}k^{2\sigma(q)}+k^{2\sigma(q)-1},$
uniformly over all $z\in\tilde{D}_{k}$ with $|z-(k+\alpha)^{2}|\gg_{A}1$ and $\|\chi\|_{L^{2q}(M)}\lesssim 1$.
Proof.
We further split $K_{\rm nr}(z)=K^{(1)}_{\rm nr}(z)+K^{(2)}_{\rm nr}(z)$, with
(128) $K^{(1)}_{\rm nr}(z)$ $=\sum_{j=1}^{d_{k}}\left(\frac{1}{\lambda_{k,j}^{2}-z}-\frac{1}{(k+\alpha)^{2}-z}\right)\chi E_{k,j}\chi,$
(129) $K^{(2)}_{\rm nr}(z)$ $=\chi\mathbf{1}(\sqrt{-\Delta_{g}}\notin[k+\alpha-1/2,k+\alpha+1/2])(-\Delta_{g}-z)^{-1}\chi,$
where $E_{k,j}$ is the orthogonal projection onto the eigenfunction of $\Delta_{g}$ corresponding to $\lambda_{k,j}^{2}$. As a consequence of (33), we have an analogous bound to (78),
(130) $\sup_{z\in\tilde{D}_{k}}\|K^{(2)}_{\rm nr}(z)\|$ $\lesssim k^{2\sigma(q)-1}.$
If we set
(131) $m_{k}(\tau,z):=\left(\frac{1}{\tau^{2}-z}-\frac{1}{(k+\alpha)^{2}-z}\right)\mathbf{1}(\tau\in[k+\alpha-A/k,k+\alpha+A/k]),$
then we have
(132) $\|K^{(1)}_{\rm nr}(z)\|$ $=\|\chi m_{k}(\sqrt{-\Delta_{g}},z)\chi\|$
(133) $\leq\|\chi\|^{2}_{L^{2q}(M)}\|m_{k}(\sqrt{-\Delta_{g}},z)\|_{L^{p}(M)\to L^{p^{\prime}}(M)}$
(134) $\lesssim\||m_{k}(\sqrt{-\Delta_{g}},z)|^{1/2}\|_{L^{p}(M)\to L^{2}(M)}^{2}.$
By $L^{2}$ orthogonality, for $z\in\tilde{D}_{k}$, $|z-(k+\alpha)^{2}|\gg A$,
(135) $\||m_{k}(\sqrt{-\Delta_{g}},z)|^{1/2}f\|_{L^{2}(M)}^{2}=\sum_{j=1}^{d_{k}}\left|\frac{1}{\lambda_{k,j}^{2}-z}-\frac{1}{(k+\alpha)^{2}-z}\right|\|E_{k,j}f\|_{L^{2}(M)}^{2}$
(136) $\leq\max_{1\leq j\leq d_{k}}\left|\frac{1}{\lambda_{k,j}^{2}-z}-\frac{1}{(k+\alpha)^{2}-z}\right|\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[k+\alpha-A/k,k+\alpha+A/k])f\|_{L^{2}(M)}^{2}$
(137) $\lesssim|(k+\alpha)^{2}-z|^{-2}k^{2\sigma(q)}\|f\|_{L^{p}(M)}^{2},$
which implies
(138) $\|K^{(1)}_{\rm nr}(z)\|\lesssim|(k+\alpha)^{2}-z|^{-2}k^{2\sigma(q)}.$
The proof is complete. ∎
Step 2: Instead of $\chi P_{k}\chi$ in the case of $S^{n}$, we now consider the nominator in the second line of (125) and denote its eigenvalues by $a_{j}(k)$, viz.
(139) $\|\chi\mathbf{1}(\sqrt{-\Delta_{g}}\in[k+\alpha-1/2,k+\alpha+1/2])\chi\|=a_{0}(k)\geq a_{1}(k)\geq\ldots\geq a_{d_{k}}(k)\geq 0.$
We now show that Lemma 3.1 still holds in this case. Let us set $\kappa=k+\alpha$. The proof of optimality of Sogge’s spectral cluster bounds shows that
(140) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\kappa-1/2,\kappa+1/2])\|_{L^{p}(M)\to L^{2}(M)}\geq c\kappa^{-n+n/p}\left(N(\kappa+1/2)-N(\kappa-1/2)\right)^{1/2}$
for some positive constant $c>0$ if $1\leq p\leq 2$ (see the display immediately after (5.1.12’) in [22]). By the sharp Weyl law and the clustering property, we have
(141) $c_{1}:=\liminf_{\kappa\to\infty}\kappa^{-n+1}\left(N(\kappa+1/2)-N(\kappa-1/2)\right)>0.$
This implies that, for $\kappa\gg 1$,
(142) $\left(N(\kappa+1/2)-N(\kappa-1/2)\right)\geq\frac{c_{1}}{2}\kappa^{n-1}$
and thus
(143) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\kappa-1/2,\kappa+1/2])\|_{L^{p}(M)\to L^{2}(M)}\geq\frac{cc_{1}}{\sqrt{2}}\kappa^{-n+n/p+\frac{n-1}{2}}$
for $\kappa\gg 1$, or equivalently, $k\gg 1$. Since $1/q=1/p-1/p^{\prime}$, the exponent of $\kappa$ equals n/(2q)1/2n/(2q)-1/2, which is $\sigma(q)$ for $q\leq(n+1)/2$. In the case $q\geq(n+1)/2$, we appeal to [22, (5.1.13)], which for $\kappa\gg 1$ gives the lower bound
(144) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[\kappa-1/2,\kappa+1/2])\|_{L^{p}(M)\to L^{2}(M)}\geq c\kappa^{(n-1)(\frac{1}{2p}-\frac{1}{4})},\quad 1\leq p\leq 2.$
The exponent of $\kappa$ equals (n1)/(4q)(n-1)/(4q), which is $\sigma(q)$ for $q\geq(n+1)/2$. By a $TT^{*}$ argument, the combination of (143), (144) yields
(145) $\|\mathbf{1}(\sqrt{-\Delta_{g}}\in[k+\alpha-1/2,k+\alpha+1/2])\|_{L^{p}(M)\to L^{p^{\prime}}(M)}\gtrsim k^{2\sigma(q)},\quad 1\leq q\leq\infty,$
for $k\gg 1$. This is the analogue, in the case of general Zoll manifolds, of (82) on $S^{n}$. Repeating the proof of Lemma 3.1, we find a nonnegative function $\chi=\chi_{k}\in L^{2q}(M)$, $\|\chi\|_{L^{2q}(M)}=1$ such that
(146) $a_{0}(k)=\|\chi\mathbf{1}(\sqrt{-\Delta_{g}}\in[k+\alpha-1/2,k+\alpha+1/2])\chi\|_{L^{2}(M)\to L^{2}(M)}\approx k^{2\sigma(q)}.$
Step 3: Recalling that $a_{j}(k)$ now denote the eigenvalues of the operator in (146), it follows that
(147) $\operatorname{spec}(\mathcal{A}_{\rm r}(z,\kappa))=\{z_{{\rm r},j}(\kappa):=(k+\alpha)^{2}+\kappa a_{j}(k):j=0,\ldots,d_{k}\},\quad a_{0}(k)\approx k^{2\sigma(q)},$
where $\mathcal{A}_{\rm r}(z,\kappa)$ is defined analogously to (88), but now with $K_{\rm r}(z)$ from (125). Choosing $\kappa_{0}$ as in (90), we get
(148) $z_{{\rm r},0}(\kappa_{0})=(k+\alpha)^{2}+\rho{\rm e}^{\mathrm{i}\theta}k^{2\sigma(q)}.$
Since the Schatten norm bound (97) of Frank and Sabin [11] for unit size spectral clusters holds on any manifold, we can repeat Steps 1-3 in the proof of Lemma 3.3. This gives us a curve
(149) $\gamma=\partial D(z_{{\rm r},0}(\kappa_{0}),\mathcal{O}(\epsilon\rho)k^{2\sigma(q)})$
with $z_{{\rm r},0}(\kappa_{0})$ as in (148) and such that
(150) $\gamma\cap\operatorname{spec}(\mathcal{A}_{\rm r}(\cdot,\kappa_{0}))=\emptyset,\quad\gamma\cap\operatorname{spec}(\mathcal{A}(\cdot,\kappa_{0}))=\emptyset,$
where $\mathcal{A}(z,\kappa_{0})$ is defined analogously to (70).
To verify the smallness condition (93), we use Lemma 3.4, which is applicable since, for $w\in\gamma$,
(151) $|w-(k+\alpha)^{2}|\geq|z_{{\rm r},0}(\kappa_{0})-(k+\alpha)^{2}|-|z_{{\rm r},0}(\kappa_{0})-w|\gtrsim\rho(1-\mathcal{O}(\epsilon))k^{2\sigma(q)}\gg_{A}1$
if $k\gg 1$ (we recall that $\sigma(q)>0$ for $q<\infty$). Since $K_{\rm r}(w)$ is a normal operator, the analogue of (112), with $(k+\alpha)^{2}$ in place of k(k+n1)k(k+n-1), holds, and thus (c.f. (117))
$\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}\|\lesssim\epsilon^{-1}.$
This, together with (127) and the bound $|\kappa_{0}|\lesssim\rho$ (c.f. (90)), yields
(152) $\max_{w\in\gamma}\|A_{\rm r}(w,\kappa_{0})^{-1}(A(w,\kappa_{0})-A_{\rm nr}(w,\kappa_{0}))\|$
(153) $\leq|\kappa_{0}|\|A_{\rm r}(w,\kappa_{0})^{-1}\|\max_{w\in\gamma}\|K_{\rm nr}\|$
(154) $\lesssim\rho\epsilon^{-1}(\max_{w\in\gamma}|(k+\alpha)^{2}-w|^{-2}k^{2\sigma(q)}+k^{2\sigma(q)-1})$
(155) $\lesssim\rho^{-1}\epsilon^{-1}k^{-2\sigma(q)}+\rho\epsilon^{-1}k^{2\sigma(q)-1}\ll 1,$
where the last inequality holds for $k\gg 1$ if $q>n/2$ and for $\rho\ll 1$, $k^{2\sigma(q)}\rho\gg 1$ for $q=n/2$, $n\geq 3$. The latter condition is satisfied by assumption.
Thus, Lemma 3.3 is also proved for the Zoll case, and the second part of Theorem 1.3 follows in the same way as the first.
4. Scaling considerations
Consider $M=\mathbb{T}^{n}$, the $n$-dimensional flat torus, viewed as $(-1/2,1/2]^{n}\subset{\mathbb{R}}^{n}$. For $L>0$, $p\in[1,\infty)$, the space $L^{p}(L\mathbb{T}^{n})$ consists of $p$-integrable functions on $Q_{L}:=(-L/2,L/2]^{n}$ satisfying periodic boundary conditions. Extension by zero yields an embedding
(156) $L^{p}(L\mathbb{T}^{n})\subset L^{p}({\mathbb{R}}^{n}).$
Conversely, multiplication with the characteristic function of $Q_{L}$ defines a projection
$\mathbf{1}_{Q_{L}}:L^{p}({\mathbb{R}}^{n})\to L^{p}(L\mathbb{T}^{n}).$
We define the unitary transformation
(157) $U_{L}:L^{2}(\mathbb{T}^{n})\to L^{2}(L\mathbb{T}^{n}),\quad(U_{L}f)(x):=L^{-\frac{n}{2}}f(x/L).$
Then
(158) $L^{-2}U_{L}(-\Delta+V)U_{L}^{-1}=-\Delta+L^{-2}V(\cdot/L),$
which implies (9). We will consider $L^{2}V(L\cdot)$ in place of $V$, in which case (9) becomes
(159) $\operatorname{spec}_{L^{2}(L\mathbb{T}^{n})}(-\Delta+V)=L^{-2}\operatorname{spec}_{L^{2}(\mathbb{T}^{n})}(-\Delta+L^{2}V(L\cdot)).$
We assume $V\in L^{q}(LM)$ for some $q\in(n/2,\infty)$; the case $q=n/2$ can be handled similarly (under an additional smallness condition), but we will not consider it here. Theorem 1.1, and more specifically (3), implies that the spectrum of $\Delta+V$ is contained in the set
(160) $\bigcup_{k=0}^{\infty}D(L^{-2}\lambda_{k}^{2},CL^{-2}r_{k})\cup\{z\in{\mathbb{C}}:L|z|^{\frac{1}{2}}(1+L^{2}|z|)^{-\sigma(q)}\leq CL^{2-\frac{n}{q}}\|V\|_{L^{q}(L\mathbb{T}^{n})}\},$
where we used the identity $L^{2}\|V(L\cdot)\|_{L^{q}(\mathbb{T}^{n})}=L^{2-\frac{n}{q}}\|V\|_{L^{q}(L\mathbb{T}^{n})}$, and where
(161) $r_{0}=L^{2-\frac{n}{q}}\|V\|_{L^{q}(L\mathbb{T}^{n})},\quad r_{k}=\lambda_{k}^{2\sigma(q)}L^{2-\frac{n}{q}}\|V\|_{L^{q}(L\mathbb{T}^{n})},\quad k\geq 1.$
Lemma 4.1.
Let $M=\mathbb{T}^{n}$ be the flat torus. Assume that $(z_{L})_{L\in{\mathbb{N}}}\subset{\mathbb{C}}$ is a convergent sequence (as $L\to\infty$).
1. If $q\in(n/2,\infty)$ and $z_{L}\in\bigcup_{k=0}^{\infty}D(L^{-2}\lambda_{k}^{2},CL^{-2}r_{k})$, then $\lim_{L\to\infty}z_{L}\in[0,\infty)$.
2. If $q\in(n/2,(n+1)/2)$ and
(162) $L|z_{L}|^{\frac{1}{2}}(1+L^{2}|z_{L}|)^{-\sigma(q)}\leq CL^{2-\frac{n}{q}}\|V\|_{L^{q}({\mathbb{R}}^{n})},$
then $z:=\lim_{L\to\infty}z_{L}$ satisfies
(163) $|z|^{\frac{1}{2}-\sigma(q)}\leq C\|V\|_{L^{q}({\mathbb{R}}^{n})}$
with the same constant $C$ as in (162).
Proof.
1. For $\epsilon>0$ to be chosen sufficiently small, we split the union of disks in (160) into two sets (strictly speaking, two sequences of sets, depending on $L$): the first set $A_{L}$ contains all disks with $\lambda_{k}\leq L^{1+\epsilon}$, and the second set $B_{L}$ contains all disks with $\lambda_{k}>L^{1+\epsilon}$. The first set satisfies
(164) $A_{L}:=\bigcup_{\lambda_{k}\leq L^{1+\epsilon}}^{\infty}D(L^{-2}\lambda_{k}^{2},CL^{-2}r_{k})\subset[0,L^{2\epsilon}]+D(0,CL^{2(1+\epsilon)\sigma(q)-\frac{n}{q}}\|V\|_{L^{q}({\mathbb{R}}^{n})}),$
where we used the embedding (156) to identify the $L^{q}$ norm of $V$ on $L\mathbb{T}^{n}$ with that on ${\mathbb{R}}^{n}$. Since $2\sigma(q)-n/q<0$ for all $q\in(n/2,\infty)$, we may choose $\epsilon$ so small (depending on $q$) that $2(1+\epsilon)\sigma(q)-\frac{n}{q}<0$. Hence, if there exists $L_{0}>1$ such that $z_{L}\in A_{L}$ for $L\geq L_{0}$, then
(165) $\lim_{L\to\infty}z_{L}\in[0,\infty).$
The alternative is that for all $L_{0}>1$, there exists $L\geq L_{0}$ such that $z_{L}\in B_{L}$. Then we can write $z_{L}=L^{-2}(\lambda_{k}^{2}+Cr_{k}{\rm e}^{\mathrm{i}\theta})$ for some $\theta\in[0,2\pi)$ and some $k$ satisfying $\lambda_{k}>L^{1+\epsilon}$. Therefore, we have
(166) $|z_{L}|\geq L^{-2}\lambda_{k}^{2}(1-Cr_{k}\lambda_{k}^{-2})\geq L^{2\epsilon}(1-o(1)).$
This shows that $(z_{L})$ does not converge, contradicting our assumption. Hence, the alternative cannot hold, and we are done.
2. Fix $R>0$. If there is a subsequence, again denoted by $(z_{L})$, such that $|z_{L}|\leq RL^{-2}$, then $\lim_{L\to\infty}z_{L}=0$, so (163) holds trivially. Otherwise, $|z_{L}|>RL^{-2}$ for $L$ sufficiently large, and then (162) implies
(167) $|z_{L}|^{\frac{1}{2}-\sigma(q)}\leq C(1+R^{-1})^{\sigma(q)}L^{1-\frac{n}{q}+2\sigma(q)}\|V\|_{L^{q}({\mathbb{R}}^{n})}.$
Since $1-\frac{n}{q}+2\sigma(q)=0$ for $q\leq(n+1)/2$ and $R$ is arbitrarily large, the claim follows. ∎